3x=4x+5
3x-4x=5
-x=5
x=-5

Found 2 solutions by lotanez, stickkens:
Answer by lotanez(23)

3x=4x+5
(-5)+ 3x = 4x + 5 -5
(-5)+ 3x = 4x
(-5)+ 3x - 3x = 4x + (-3x)
(-5)= 1x
-5 = x or x=-5

Answer by stickkens(2)